A steel storage tank for propane gas is to be constructed in the shape 

of a right circular cylinder with a hemisphere at both ends. If the 

desired capacity is 100 cu. ft., what dimensions will require the least 

amount of steel ?


1
Expert's answer
2021-03-31T14:23:59-0400

Consider the right circular cylindrical storage tank with a hemisphere at both ends.


Let "h" be the height of the cylindrical section and "r" be the radius of cylinder.


Obviously, the radius of hemisphere will also be "r"


The total volume of the storage tank is "100 ft^3"


So,


"\pi r^2 h+2(\frac{2}{3} \pi r^3)=100"


"\pi r^2 h+\frac{4}{3} \pi r^3=100"


"3\pi r^2 h+4 \pi r^3=300"


"3\pi r^2 h=300-4 \pi r^3"


"h=\frac{300-4 \pi r^3}{3\pi r^2}"



Now, the total surface area of the tank is,


"S=2 \pi r h+4 \pi r^2"


Substitute "\frac{300-4 \pi r^3}{3\pi r^2}" for "h" into "S=2 \pi r h+4 \pi r^2" to obtain the total surface area in a single variable "r" as,


"S(r)=2 \pi r (\frac{300-4 \pi r^3}{3\pi r^2})+4 \pi r^2"


"S(r)=\frac{200}{r}-\frac{8 \pi}{3}r^2+4 \pi r^2"


"S(r)=\frac{200}{r}+\frac{4 \pi}{3}r^2"


Differentiate "S" with respect to "r" and set it equal to zero as,


"S'(r)=0"


"-\frac{200}{r^2}+\frac{8 \pi}{3}r=0"


"\frac{8 \pi}{3}r=\frac{200}{r^2}"


"r^3=\frac{75}{\pi}"


"r=\sqrt[3]{\frac{75}{\pi}}"


Here, "S"(r)=\frac{400}{r^3}+\frac{8 \pi}{3}>0," therefore, "S" is minimum.


Plug "r=\sqrt[3]{\frac{75}{\pi}}" into "h=\frac{300-4 \pi r^3}{3\pi r^2}" and simplify for "h" as,


"h=\frac{300-4 \pi (\frac{75}{\pi})}{3\pi (\sqrt[3]{\frac{75}{\pi}})^2}=0"


Therefore, the dimensions are: "r=\sqrt[3]{\frac{75}{\pi}}" and "h=0" .